OpenStudy (johnw):

f(x) = arccos(x^2+1) What is the derivative of the function?

OpenStudy (kainui):

Implicit differentiation is the best choice here. Take cosine of both sides to get: \[\cos(f)=x^2+1\] Then differentiate with respect to x. \[-\sin (f) * f' = 2x\] Now you can figure out what \(\sin f\) is from \(\cos f\) by remembering the pythagorean identity: \[\cos^2 f + \sin^2 f = 1\]\[\sin f = \sqrt{1-\cos^2 f} = \sqrt{1-(x^2+1)^2}\] Plug in and you're done!

OpenStudy (johnw):

I´m not sure I follow your answer

OpenStudy (johnw):

This is really not my area of expertise

OpenStudy (davejavous):

First of all, lets set the question up: \[f(x) = acos(x^2+1) \]\[let \; t=x^2+1 \]\[\therefore\ f(x) =acos(t)\] The rule for differentiating acos is: \[{\frac{ d }{ dx } \left[acos(x)\right] = \frac{ -1 }{ \sqrt{1-x^2}}}\] But this is a chain rule question. The chain rule is: \[\frac{df}{dx} = \frac{df}{dt} \times \frac{dt}{dx} \] Substituting into the chain rule gives: \[ \frac{df}{dx} = \frac{ -1 }{ \sqrt{1-t^2}} \times 2x \] Substituting t into the equation and also simplifying the answer gives us: \[ \frac{df}{dt} = \frac{2x}{\sqrt{1 - (x^2 + 1)^2}} \]

OpenStudy (johnw):

Could you try explaining using the chain rule (f o g)´(x)=f´(g(x))*g´(x) if it´s not too much trouble?

OpenStudy (johnw):

I really appreciate your help, but I´ve been using this instead. Well really it´s just (f o g)´(x) right?

OpenStudy (davejavous):

The two functions are: \[f(x) = acos(x) \]\[ g(x) = x^2 + 1\] The differentials are: \[f'(x) = \frac{-1}{\sqrt{1-x^2}}\]\[g'(x) = 2x\] Using the chain rule: \[ (f \circ g)'(x) = f'(g(x)) \times g'(x) \] Then substituting in f and g: \[ (f \circ g)'(x) = \frac{-1}{\sqrt{1-(x^2+1)^2}} \times 2x\] Therefore the final answer is: \[ (f \circ g)'(x) = \frac{-2x}{\sqrt{1-(x^2+1)^2}}\]

OpenStudy (johnw):

Thank you so much!

OpenStudy (kainui):

There seems to be a fatal flaw here that really bothered me last night.

OpenStudy (kainui):

Specifically when we get to the end, we have the square root of a negative number. So I got to thinking, OK what's the domain of \(\arccos\)? It's \([-1,1]\). What's the range of \(x^2+1\)? It's \([1,\infty)\). So the only point this "function" works for is \(x=0\).

OpenStudy (johnw):

I believe you're right. But the derivative of arccos(x^2+1) is still the earlier mentioned answer? Was the domain really the issue here?

OpenStudy (kainui):

It's the derivative of nothing though. The answer comes out to be nonsense... \[f'=\frac{-2}{\sqrt{-x^2-2}}\]

OpenStudy (naka354):

f'(x) = (-2x)/√(-x^4 - 2x^2)